The difference of the square of a number and 4 is equal to 3 times that number. Find the negative solution.

Answer by fractalier(6550)

Call the number, x. Then we write
x^2 - 4 = 3x
Now solve for the negative solution for x...
x^2 - 3x - 4 = 0
Factoring
(x - 4)(x + 1) = 0
x = 4 and x = -1
So x = -1.